OpenStudy (anonymous):

Solve the linear system by using substitution. 6x-y=-35 5x-2y=-35 I've done this twice already and both times I've come out with x=15 and y=125. I've plugged those into the original equations but it doesn't work on the second equation. Help?

OpenStudy (anonymous):

\[6x-y=-35\] \[y=6x+35\] \[5x-2(6x+35)=-35\] \[5x-12x-70=-35\] \[-7x=35\] \[x=-5\] \[5(-5)-2y=-35\] \[-10=-2y\] \[5=y\]

OpenStudy (anonymous):

hopefully you understand this work seen as though it sounds as though you've answered some of these problems before . If you need help let me know ;)

OpenStudy (anonymous):

Oh thanks, I see what I did. In your third step, you multiplied by negative 2 while I did it with a negative. Thanks for your help. ^_^